OpenStudy (anonymous):

Having trouble setting these problems up :/ a) Mrs.Sheik was 24 years old when Ingrid was born. In three years, the sum of their ages will be 68. how old is each now? b) Gunta's dad is 25 years older than Gunta. The sum of their ages will be 61. How old is Gunta?

OpenStudy (ash2326):

a) Let Sheik and Ingrid's age be x and y, When Ingrid was born Sheik was 24 so we get \[x-y=24\] In three years the sum of their ages is 68, so \[(x+3)+(y+3)=68\] I think you can do now, @GOODMAN

OpenStudy (anonymous):

Ohhhh..okayy!!! Thanks! The same applies to the other problem as well?

OpenStudy (ash2326):

Yeah, Could you write the equation for second one?

OpenStudy (anonymous):

Let Gunta=x and Dad=y Gunta's dad is 25 years older than Gunta, so, y=x+25 The sum of their ages is 61 x+y=61 I dont know if i did it right...

OpenStudy (ash2326):

Correct:D Good work @GOODMAN

OpenStudy (anonymous):

So exactly how would i find the ages..im still a little confused about all the variables.

OpenStudy (ash2326):

In both the cases, x and y will give us the present ages

OpenStudy (anonymous):

Okay!

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

Yea, somewhat! I got the idea, im just not fond of word problems, lol. Thanks for your help!

OpenStudy (ash2326):

You're welcome, :D

OpenStudy (anonymous):

@OpenstudyUser